OpenStudy (anonymous):

solve each system using substitution. check your answer. y=8-x 7=2-y

OpenStudy (dtan5457):

y=8-x 7=2-y if y=8-x, plug that into 7=2-y 7=2-(8-x)

OpenStudy (anonymous):

what do i do then

OpenStudy (dtan5457):

Solve for x..

OpenStudy (anonymous):

would it be 7=16-x

OpenStudy (anonymous):

hello

OpenStudy (dtan5457):

i just noticed something " 7=2-y" where's the x?